HelpMePlz:

Which ordered pair is the solution to the system of equations? 2x+y=-8 x-y=-4 A. (−3, −2) B. (−1, 3) C. (1, 5) D. (−4, 0)

Champion:

I don't recommend short cuts for multiple choice questions, because real problems in life are seldom multiple choice questions. However, for this particular type of problem where we're given multiple answers to a system of equations, instead of checking one equation at a time, I would add the equations together and in this case, I would check the sum, equation (3) against the answer choices. 2x+y=-8...(1) x-y=-4......(2) (1)+(2): 3x=-12 ........(3) or x=-4 There is only one answer choice that gives x=-4, so I would go for it, and check that answer choice against each of equations (1) and (2). This is especially time-saving for the systems of more than two variables.